OpenStudy (anonymous):

A particular reactant decomposes with a half-life of 175 s when its initial concentration is 0.375 M. The same reactant decomposes with a half-life of 205 s when its initial concentration is 0.320 M. What will be the Rate constant?

OpenStudy (anonymous):

It's a second order reaction!

OpenStudy (jfraser):

You know the order, and you know that the rate will be some change in concentration over time. The half-life info gives you a way to find rate. Use those values to find k.

OpenStudy (anonymous):

Nt=Noe^-kt T1/2=In2/K so i think these are the formulas we ll have to figure out how to go from here

OpenStudy (aaronq):

^ thats for a first order, the half life equation for a second order would be: \[t _{1/2}=\frac{ 1 }{ k[A]_{0} }\]